Attendance at a university’s basketball games follows a normal distribution with mean
μ = 8,000 and standard deviation σ = 1,000. Estimate the percentage of games that
have between 7,000 to 9,000 people in attendance.
Answer #1


P ( 7000 < X < 9000 )
Standardizing the value

Z = ( 7000 - 8000 ) / 1000
Z = -1
Z = ( 9000 - 8000 ) / 1000
Z = 1
P ( -1 < Z < 1 )
P ( 7000 < X < 9000 ) = P ( Z < 1 ) - P ( Z < -1 )
P ( 7000 < X < 9000 ) = 0.8413 - 0.1587
P ( 7000 < X < 9000 ) = 0.6827

0.6827 * 100 = 68.27%  of games that have between 7,000 to 9,000 people in attendance.